OpenStudy (anonymous):

Choose the ratio below that is equivalent to 3 : 5. A. 8 : 4 B. 9 : 15 C. 5 : 3 D. 9 : 5

OpenStudy (anonymous):

it would be d

OpenStudy (texaschic101):

wrong

OpenStudy (anonymous):

why its right

OpenStudy (texaschic101):

what is equivalent to 3:5 which equals 3/5 8:4 = 8/4 9:15 = 9/15 = 3/5 5:3 = 5/3 9:5 = 9/5

OpenStudy (anonymous):

ok but shes asking which one not how many

OpenStudy (texaschic101):

The equivalent one is 9:15 which equals 9/15 which equals 3/5 which equals 3:15.

OpenStudy (texaschic101):

oops 3/5 = 3:5